(a) John, aged 70, has a house worth $500,000. He intends to borrow $100,000 as a lump sum. If the interest rate is 10% p.a. payable annually (i=10%), and house prices increase by 3% p.a. (f = 3%), what is the percentage equity in 15 years?
Answer #1

Current worth of house=$500,000

Annual increase of house price=3%=0.03

Value of house after 15 years=500000*(1.03^15)=$778,984

Interest rate on loan=10%

Amount of Loan =$100,000

Loan Balance after 15 years=100000*(1.1^15)=$417,725

Home Equity =778984-417725=$361,259

Percentage of Equity =(361259/778984)*100%=46.38%